Hero Manufacturing has 7.4 million shares of common stock outstanding. The current share price is $88 and the book value per share is $3. The company also has two bond issues outstanding. The first bond issue has a face value of $65 million, a coupon rate of 6.2 percent and sells for 106.3 percent of par. The second issue has a face value of $45.5 million, a coupon rate of 6.7 percent and sells for 110.1 percent of par. The first issue matures in 7 years, the second in 23 years.

Suppose the company’s stock has a beta of 1.3. The risk-free rate is 2.7 percent and the market risk premium is 6.6 percent. Assume that the overall cost of debt is the weighted average implied by the two outstanding debt issues. Both bonds make semiannual payments. The tax rate is 21 percent. What is the company’s WACC? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

Number of Shares Outstanding = 7.4 million, Current Stock Price = $ 88, Book Value per Share = $ 3,

Market Value of Equity = E = 7.4 x 88 = $ 651.2 million

Risk-Free Rate = 2.7%, Market Risk Premium = 6.6 %, Beta = 1.3

Cost of Equity = 2.7 + 1.3 x 6.6 = 11.28 %

Debt:

Bond 1: Face Value = $ 65 million, Coupon Rate = 6.2 %, Price = 106.3 % of Par Value = 1.063 x 65 = $ 69.095 million, Tenure = 7 years

Coupon = 65 x 0.062 = $ 4.03 million

Let the bond yield be y1

Therefore, 60.095 = 4.03 x (1/y1) x [1-{1/(1+y1)^(7)}] + 69.095 / (1+y1)^(7)

Using EXCEL's Goal Seek Function to solve the above equation, we get:

y1 = 0.0510707 or 5.10707 % ~ 5.107%

Bond2: Face Value = $ 45.5 million, Tenure = 23 years, Interest Rate = 6.7 % and Price = 110.1 % of Face Value = 1.101 x 45.5 = $ 50.0955 million

Coupon = 45.5 x 0.067 = $ 3.0485 million

Let the bond yield be y2

Therefore, 50.0955 = 3.0485 x (1/y2) x [1-{1/(1+y2)^(23)}] + [45.5/(1+y2)^(23)]

Using EXCEL's Goal Seek Function to solve the above equation, we get:

y2 = 0.058874 or 5.8874 % ~ 5.89 %

Total Debt Value = 50.0955 + 69.095 = $ 119.1905 million

Total Firm Value = 651.2 + 119.1905 = $ 770.3905 million

Equity Proportion = (651.2/770.3905) = 0.8453 and Debt Proportion = 1-0.8453 = 0.1547

Weighted Average COst of Debt = (50.0955/119.1905) x 5.89 + (69.095/119.1905) x 5.107 = 0.052636 or 5.2636 % ~ 5.26%

Tax Rate = 21 %

Weighted Average Cost of Capital = 0.8453 + 11.28 + 0.1547 x 5.26 x (1-0.21) = 10.1778 % ~ 10.18 %